OpenStudy (lexy148):

what is 2(2/3) X 1/7 ?

OpenStudy (skullpatrol):

any ideas?

OpenStudy (skullpatrol):

$$\Huge \dfrac{a}{b}\times\dfrac{c}{d}=\dfrac{a\times c}{b\times d}$$